OpenStudy (anonymous):

3/4x+4 = 5/8 solve for x

OpenStudy (ksaimouli):

3/4x=5/8-4

OpenStudy (ksaimouli):

(4(5/8 -4))/3

OpenStudy (anonymous):

x=−9/2

OpenStudy (anonymous):

I would start by changing 3/4x to 6/8x Then multiply everything by 8 to get 6x+32=5 Subtract 32 on both sides. 6x=-27 Divide by 6 on both sides x=-27/6 Simplify by dividing the top and bottom by 3 x=-9/2 Just like @david111 said.

OpenStudy (anonymous):

thanks i got that to but -9/2 isnt in the choices

OpenStudy (anonymous):

Strange. What are the choices? And your 100% that you typed the question incorrectly?

OpenStudy (anonymous):

a. -4 1/8 b. -4 1/2 c. 3 1/3 d. 6 1/6

OpenStudy (anonymous):

nvm i got it its b

OpenStudy (anonymous):

Yupp!